What is the ratio of two wires having same length, but diameter of first wire is just doubled that of second wire?

Grade:12th pass

1 Answers

Khimraj
6 years ago
I think youare asking about ratio of resistance
as R \propto l/A \propto l/d2
where l is length and d is diameter of wire.
Then
R1/R2 = (d2/d1)2 = (1/2)2 = 1/4